OpenStudy (goldphenoix):

What does 0^0 give us? Explain your answer please.

OpenStudy (valpey):

We get to define it for our purposes. It is usually best to let the rule "anything to the zero power equals one" rule trump the "zero to any power is zero" rule because of some cool continuity phenomena of various sorts. I am thinking of the Gamma function or something.

OpenStudy (goldphenoix):

So your claiming that 0 to the 0th power is zero?

OpenStudy (valpey):

No, 1. It makes things like: \[e^x = \sum_{n=0}^{\infty}\frac{x^n}{n!}\]work even for x = 0.

OpenStudy (goldphenoix):

I'm going to 8th grade. I do not understand this. :|

OpenStudy (valpey):

0^0 = 1. But only really because that is how we prefer to define it. It makes writing a lot of cool equations simpler.

hero (hero):

0^0 violates rules of exponents because they can't figure out if it equals zero or one, so they just say it is undefined.

OpenStudy (goldphenoix):

What if 0^0 is in a problem? Do I consider it as 1 or 0?

OpenStudy (valpey):

Consider it as 1.

OpenStudy (whpalmer4):

No, 0^0 is indeterminate. It is sometimes defined to equal 1 for convenience.

OpenStudy (goldphenoix):

Ah, thanks. Hardest part. Who do I give the medal to? =/

OpenStudy (whpalmer4):

I'll make that part slightly easier — don't give it to me.

OpenStudy (goldphenoix):

Thank you everyone! :)

hero (hero):

If you tell someone not to do something...

OpenStudy (goldphenoix):

What do you mean, Hero?

OpenStudy (valpey):

http://en.wikipedia.org/wiki/Exponential_function The expansion only works if e^0 = 1. Which is only true if 0^0 = 1.